What is the resistance and temperature coefficient of resistance of a resistor if its color code is yellow-green-blue-violet-red-yellow?



1
Expert's answer
2022-04-01T16:32:31-0400

Resistance

V=iR

"R=R_0(1+\alpha(T-T_0))"


"R=45\times10^6+45\times10^6\times\frac{5}{100}=45\times10^{6}[1+0.5]"

"R=72\times10^4+72\times10^4\times\frac{5}{100}=72\times10^{4}[1+0.5]"